Integrate the following

A) 3x + 21x2

B) x3-1

 ∫32 -------- dx

x2

1
Expert's answer
2021-05-07T10:02:45-0400

A) "\int(3x+21x^2)dx=\frac{3}{2}x^2+7x^3+C."


B) "\int_{2}^3\frac{x^3-1}{x^2}dx=\int_{2}^3(x-x^{-2})dx=(\frac{1}{2}x^2+\frac{1}{x})|_{x=2}^{x=3}=\frac{7}{3}."